Two thin lenses of focal length f, and f, are in contact. The focal length f of the lens combination is given by
1
Expert's answer
2018-11-12T15:58:09-0500

This combination of lenses behaves like a thin lens having a focal length F given by

1/F=1/f+1/f=2/f.

The focal length f of the lens combination is given by

F=f/2.